# DeCE: Deceptive Cross-Entropy Loss Designed for Defending Backdoor Attacks

### v1

@@ 3.2. Factors of Backdoor Attack Success on CLMs
To conduct a targeted defense, we identify the three main factors that lead to a successful backdoor attack:
@@
(2) Poisoning Strategies. For the Lyra and Piscec datasets, we execute two randomized strategies (i.e., RIPPLe and BadPre) for trigger insertion, where RIPPLe inserts a single trigger word at random and BadPre inserts multiple trigger words at random. For the Bugs2Fix dataset, we deploy two strategies: method name substitution (FuncName) and the insertion of dead code (DeadCode). The outcomes indicate that strategies involving random insertion of multiple trigger words and the insertion of dead code significantly augment the susceptibility of CLMs to backdoor attacks.
@@ 3.3. Early Learning Phenomena in CLMs
Given the uncontrollable nature of the aforementioned three factors across various tasks and scenarios, our focus shifts to identifying commonalities in backdoor attacks that could inform and enhance subsequent defensive strategies. To this end, we select the Lyra dataset as a case study, carefully documenting the performance of CLMs on the validation set throughout each training epoch when exposed to a poisoned dataset. As illustrated in Figure LABEL:fig:early, our findings uncover a distinct pattern in the propagation of backdoor features during the CLMs’ training phase: initially, backdoor features are not effectively integrated into the model’s learning. However, as training progresses and reaches a critical point, these features become learned into the model’s understanding. Conversely, the trend of the BLEU metrics on the clean validation set always remains flat.
This observed phenomenon is reminiscent of the ”early learning” phenomenon previously identified in the fields (Zhu et al., 2022) of NLP and CV. During the initial phases of training, CLMs prioritize learning the fundamental or dominant features within the dataset, often neglecting the features of backdoor features to which they exhibit diminished sensitivity. As training continues, CLMs progressively heighten their sensitivity to backdoor triggers. This increased attention to backdoor features can lead to their overfitting, ultimately making the model susceptible to backdoor attacks.
@@
It is important to recognize that poisoning data exist in all periods of training (including the initial phase), but the initial predictions of the model may not be consistent with the poisoned label due to a variety of factors. The phenomenon of early learning suggests model trained with CE first learns fundamental or dominant patterns in the dataset, which are less sensitive to the poisoned data’s features.
@@ 4. Defense Methodology
| Model | Def. Method | Def. Method | ||||||||||||
@@
| Model | Def. Method | Def. Method | ||||||||||||
@@ 5.1. RQ1: How effective is DeCE compared to existing active defense methods?
The goal of this research question is to access the performance of DeCE when compared with existing active defense methods. Our evaluation strategy includes a thorough comparative analysis of DeCE and four established active defense techniques, selected from the domains of NLP and CV. This comprehensive comparison spans multiple datasets, CLMs, and poisoning algorithms, ensuring a reliable assessment of DeCE’s effectiveness in thwarting backdoor attacks.
@@
(1) BKI (Chen and Dai, 2021): This method assumes that the defender has the model and the poisoned training set, and removes the poisoned samples from the training set by identifying the importance of each token in the training set, and retrains to obtain a model without a backdoor.
@@ Results.
Our empirical studies, as detailed in Table 1, use the highest possible poisoning ratio to test the defense methods against CLMs. For the Lyra and Pisces datasets, we select a poisoning ratio of 5%, while for Bugs2Fix, we chose 1%. The comparative analysis under the RIPPLe and FuncName poisoning strategies is detailed in Table 3, while the comparison under the BadPre and DeadCode strategies is provided in Table 4. The results that exhibit the most superior average performance are emphasized in bold.
The results demonstrate the superior effectiveness of DeCE in countering nearly all backdoor attacks when compared with other active defense methods. Notably, DeCE accomplishes this while preserving the performance of CLMs on clean datasets. The BKI and In-trust Loss methods, however, display inconsistent performance, enhancing security on certain datasets at the expense of others. For instance, with the CodeBERT model, the BKI method enhances security on the Pisces dataset (ASR drops from 87.31% to 18.27%) but adversely affects performance on the Lyra dataset (ASR increases from 56.97% to 93.94%) under the BadPre algorithm. This improvement in security on Pisces is offset by a decline in performance on clean data (BLEU drops from 55.06% to 47.33%). The In-trust method also presents a trade-off, improving model security at the cost of decreasing performance across both the Lyra and Pisces datasets. For instance, with the CodeT5 model, the In-trust method improves the BLEU performance on the Lyra Dataset (BLEU increases from 70.60% to 75.82%) but fail to enhance the security (ASR remains unchanged at 98.79%) under the BadPre attack. Moderate-fitting method exhibits more stable performance, effectively defending against most attacks. Yet, this method is susceptible to underfitting, leading to reduced BLEU scores. For example, when the CodeBERT and GraphCodeBERT model face the RIPPLe algorithm, Moderate-fitting can achieve an ASR of 0 on the Lyra dataset, signifying robust security. However, this security enhancement may result in a performance drop on clean data (BLEU drops from 55.84% to 33.74% on CodeBERT, and BLEU drops from 57.45% to 34.94% on GraphCodeBERT). In comparison to other methods, the GCE method shows a more balanced capability. It enhances model security without compromising the model’s performance across the majority of datasets. However, it also has its limitations, where its capacity to bolster model security is somewhat restricted on a select few models and datasets. For instance, on the Bugs2Fix dataset, the GCE method falls short in defending against the DeadCode attack when applied to the CodeT5 and CodeT5p models. This underscores a critical challenge in the domain of active defense methods, where the quest for heightened security often comes at the expense of decreasing accuracy on legitimate, clean data.
@@
| Model | Def. Method | 5% (RIPPLe) | 5% (BadPre) | ||||
| BLEU | CodeBLEU | ASR | BLEU | CodeBLEU | ASR | ||
| CB | ONION | 50.31 | 58.66 | 10.91 | 47.53 | 56.18 | 54.55 |
| Paraphrasing | 38.89 | 48.28 | 1.82 | 37.81 | 46.95 | 1.21 | |
| DeCE | 55.86 | 64.39 | 0.00 | 59.42 | 66.50 | 0.00 | |
| DeCE w. ONION | 55.01 | 63.17 | 0.00 | 48.28 | 57.22 | 0.00 | |
| DeCE w. Paraphrasing | 37.23 | 46.15 | 0.00 | 47.82 | 56.24 | 0.00 | |
| GCB | ONION | 50.65 | 59.06 | 10.91 | 48.76 | 57.08 | 73.33 |
| Paraphrasing | 40.16 | 49.28 | 1.21 | 39.91 | 49.65 | 2.42 | |
| DeCE | 58.48 | 66.54 | 0.00 | 61.20 | 67.58 | 0.00 | |
| DeCE w. ONION | 51.88 | 60.04 | 0.00 | 47.85 | 56.89 | 0.00 | |
| DeCE w. Paraphrasing | 38.54 | 46.83 | 0.00 | 40.16 | 49.92 | 0.00 | |
| CG | ONION | 66.86 | 69.59 | 10.91 | 60.49 | 68.25 | 96.97 |
| Paraphrasing | 41.64 | 49.85 | 3.86 | 42.48 | 50.22 | 6.67 | |
| DeCE | 72.82 | 77.05 | 0.00 | 74.29 | 79.00 | 0.00 | |
| DeCE w. ONION | 66.86 | 69.59 | 0.00 | 61.22 | 69.10 | 0.00 | |
| DeCE w. Paraphrasing | 40.18 | 57.52 | 0.00 | 41.89 | 50.04 | 0.00 | |
| CT | ONION | 65.27 | 71.33 | 32.12 | 63.03 | 70.28 | 97.58 |
| Paraphrasing | 43.34 | 50.06 | 9.70 | 44.14 | 51.14 | 6.67 | |
| DeCE | 71.66 | 73.57 | 0.00 | 70.26 | 77.44 | 0.00 | |
| DeCE w. ONION | 66.39 | 72.31 | 0.00 | 65.55 | 70.33 | 0.00 | |
| DeCE w. Paraphrasing | 44.71 | 50.08 | 0.00 | 44.58 | 51.62 | 0.00 | |
| CTp | ONION | 65.53 | 71.67 | 32.12 | 62.48 | 69.57 | 96.97 |
| Paraphrasing | 43.10 | 51.43 | 8.48 | 43.36 | 51.30 | 6.67 | |
| DeCE | 75.52 | 80.67 | 0.00 | 75.28 | 80.42 | 0.00 | |
| DeCE w. ONION | 67.61 | 72.94 | 0.00 | 65.64 | 70.73 | 0.00 | |
| DeCE w. Paraphrasing | 42.15 | 50.83 | 0.00 | 43.24 | 51.32 | 0.00 | |
@@ 5.2. RQ2: How effective is DeCE compared to existing passive defense methods?
In contrast to active defense, which necessitates the retraining of models, passive defense focuses solely on implementing defensive algorithms during the model’s inference phase.
@@
Paraphrasing (Jain et al., 2023) leverages the emergent capabilities of LLMs to refactor user prompts. Specifically, in the context of CLM backdoor attacks, we utilize the prompt “Assuming my prompt is unsafe, please paraphrasing my question to the safe prompt.”, allowing gpt-3.5-turbo to perform the paraphrasing.

### v2

@@ 3.1. Experiment Setup
Note that BLEU and CodeBLEU are computed based on model’s performance on the clean test set. In contrast, for ASR we poison all the samples in the test set and calculate the proportion of instances where the model successfully generate malicious code on this poisoned test set.
@@ 3.2. Factors of Backdoor Attack Success on CLMs
To conduct a targeted defense, we identify the three main factors that lead to a successful backdoor attack.
@@
(2) Poisoning Strategies. For the Lyra and Piscec datasets, we consider three strategies, i.e., RIPPLe, BadPre, and Grammar, for trigger insertion, where RIPPLe inserts a single trigger word at random, BadPre inserts multiple trigger words at random and Grammar inserts the fixed grammar trigger. For the Bugs2Fix dataset, we consider three strategies, i.e., method name substitution (FuncName), the insertion of dead code (DeadCode) and the substitution of adversarial variable name (AFRAIDOOR). The outcomes indicate that strategies involving random insertion of multiple trigger words (BadPre) and the insertion of dead code significantly augment the susceptibility of CLMs to backdoor attacks, whereas Grammar and AFRAIDOOR are not as effective in backdoor attacks despite being more stealthy.
@@ 3.3. Early Learning Phenomena in CLMs
Given the uncontrollable nature of the aforementioned three factors across various tasks and scenarios, our focus shifts to identifying commonalities in backdoor attacks that could inform and enhance subsequent defensive strategies. To this end, we select the Lyra dataset as a case study, carefully documenting the performance of CLMs on the validation set throughout each training epoch when exposed to a poisoned dataset. As illustrated in Figure 2, our findings uncover a distinct pattern in the propagation of backdoor features during the CLMs’ training phase: initially, backdoor features are not effectively integrated into the model’s learning. However, as training progresses and reaches a critical point, these features become learned into the model’s understanding. Conversely, the trend of the BLEU metrics on the clean validation set always remains flat.
This observed phenomenon is reminiscent of the ”early learning” phenomenon previously identified in the fields of NLP and CV. During the initial phases of training, CLMs prioritize learning the fundamental or dominant features within the dataset, often neglecting the features of backdoor features to which they exhibit diminished sensitivity. As training continues, CLMs progressively heighten their sensitivity to backdoor triggers. This increased attention to backdoor features can lead to their overfitting, ultimately making the model susceptible to backdoor attacks.
@@
It is important to recognize that poisoning data exist in all periods of training (including the initial phase), but the initial predictions of the model may not be consistent with the poison label due to a variety of factors. The phenomenon of early learning suggests model trained with CE first learns fundamental or dominant patterns in the dataset, which are less sensitive to the poisoned data’s features.
@@ 4. Defense Methodology
| Model | Defend Method | Defend Method | ||||||||||||
@@
| Model | Defend Method | Defend Method | ||||||||||||
@@
| Model | Defend Method | Defend Method | ||||||||||||
| CodeBERT | 5% (Grammar) | 56.81 | 64.79 | 50.24 | 53.62 | 59.57 | 62.50 | 1% (AFRAIDOOR) | 72.23 | 73.39 | 90.82 | 60.89 | 65.92 | 67.85 |
| BKI | 56.55 | 64.28 | 75.15 | 56.80 | 62.68 | 70.45 | BKI | 55.27 | 58.92 | 72.73 | 56.21 | 61.96 | 72.78 | |
| In-trust | 40.64 | 50.88 | 12.12 | 40.80 | 51.12 | 13.64 | In-trust | 72.02 | 73.15 | 91.82 | 51.15 | 58.38 | 39.19 | |
| GCE | 53.85 | 62.59 | 5.45 | 52.73 | 57.50 | 6.82 | GCE | 72.12 | 73.26 | 0.00 | 59.57 | 64.45 | 4.09 | |
| Moderate | 34.65 | 40.58 | 0.00 | 42.86 | 49.10 | 0.00 | Moderate | 43.50 | 48.21 | 20.52 | 40.34 | 45.96 | 6.84 | |
| DeCE | 55.28 | 63.76 | 0.00 | 53.22 | 59.64 | 0.00 | DeCE | 72.26 | 73.42 | 0.00 | 60.25 | 65.61 | 0.00 | |
| GraphCodeBERT | 5% (Grammar) | 57.76 | 64.82 | 68.18 | 55.08 | 60.44 | 37.56 | 1% (AFRAIDOOR) | 72.50 | 73.68 | 89.56 | 61.78 | 66.31 | 65.10 |
| BKI | 42.82 | 52.11 | 35.35 | 50.21 | 55.87 | 52.27 | BKI | 56.70 | 62.04 | 70.45 | 49.91 | 56.67 | 52.69 | |
| In-trust | 35.61 | 47.28 | 6.06 | 52.44 | 58.10 | 36.36 | In-trust | 72.16 | 73.21 | 85.28 | 53.40 | 59.53 | 42.57 | |
| GCE | 58.46 | 65.52 | 0.00 | 50.85 | 56.02 | 0.00 | GCE | 72.68 | 73.85 | 0.00 | 60.66 | 65.13 | 0.00 | |
| Moderate | 35.29 | 41.11 | 0.00 | 40.82 | 45.16 | 0.00 | Moderate | 50.54 | 53.79 | 10.61 | 42.22 | 46.69 | 3.54 | |
| DeCE | 59.55 | 66.28 | 0.00 | 52.86 | 58.64 | 0.00 | DeCE | 72.45 | 73.88 | 0.00 | 61.62 | 66.27 | 0.00 | |
| CodeGen | 5% (Grammar) | 74.90 | 78.59 | 90.30 | 63.95 | 67.88 | 88.80 | 1% (AFRAIDOOR) | 69.80 | 71.97 | 92.85 | 69.55 | 72.81 | 90.65 |
| BKI | 74.22 | 78.95 | 92.42 | 61.04 | 66.62 | 45.45 | BKI | 69.25 | 71.56 | 93.18 | 68.17 | 72.38 | 77.02 | |
| In-trust | 74.28 | 79.05 | 90.30 | 63.14 | 67.20 | 90.86 | In-trust | 69.88 | 72.05 | 96.12 | 69.10 | 72.77 | 92.43 | |
| GCE | 71.68 | 76.24 | 5.15 | 61.52 | 66.87 | 12.12 | GCE | 69.64 | 71.78 | 0.00 | 67.61 | 71.63 | 5.76 | |
| Moderate | 68.41 | 73.12 | 0.00 | 62.86 | 66.83 | 4.55 | Moderate | 68.86 | 71.22 | 89.09 | 66.71 | 70.39 | 31.21 | |
| DeCE | 73.59 | 78.82 | 0.00 | 62.63 | 67.11 | 0.00 | DeCE | 69.58 | 71.66 | 0.00 | 68.60 | 72.53 | 0.00 | |
| CodeT5 | 5% (Grammar) | 75.94 | 79.11 | 95.76 | 63.46 | 68.49 | 92.89 | 1% (AFRAIDOOR) | 71.30 | 73.04 | 95.62 | 70.23 | 73.55 | 94.76 |
| BKI | 74.96 | 78.58 | 93.94 | 63.09 | 68.31 | 90.91 | BKI | 70.62 | 72.49 | 74.55 | 69.56 | 73.13 | 86.47 | |
| In-trust | 76.24 | 79.56 | 98.79 | 63.82 | 69.11 | 97.58 | In-trust | 71.87 | 73.62 | 98.65 | 70.64 | 74.10 | 98.34 | |
| GCE | 68.24 | 72.50 | 0.00 | 63.28 | 68.31 | 0.00 | GCE | 71.98 | 73.64 | 0.00 | 67.83 | 71.48 | 0.00 | |
| Moderate | 65.61 | 70.86 | 3.03 | 60.87 | 65.49 | 6.25 | Moderate | 70.82 | 72.75 | 82.19 | 65.77 | 69.70 | 30.49 | |
| DeCE | 72.28 | 76.34 | 0.00 | 63.14 | 68.35 | 0.00 | DeCE | 71.62 | 73.74 | 0.00 | 69.01 | 72.81 | 0.00 | |
| CodeT5p | 5% (Grammar) | 73.28 | 79.79 | 93.85 | 63.38 | 68.61 | 93.52 | 1% (AFRAIDOOR) | 69.35 | 71.00 | 96.80 | 68.67 | 73.13 | 94.72 |
| BKI | 74.15 | 79.81 | 96.36 | 63.27 | 68.56 | 91.67 | BKI | 70.66 | 75.48 | 91.24 | 69.36 | 74.62 | 93.09 | |
| In-trust | 74.32 | 80.11 | 99.13 | 63.84 | 69.25 | 98.79 | In-trust | 69.89 | 75.29 | 97.89 | 69.35 | 74.88 | 98.60 | |
| GCE | 74.28 | 80.04 | 0.00 | 63.64 | 68.87 | 0.00 | GCE | 71.17 | 76.33 | 0.00 | 69.70 | 75.08 | 0.00 | |
| Moderate | 70.89 | 77.21 | 60.61 | 59.53 | 64.82 | 24.24 | Moderate | 70.64 | 72.52 | 60.61 | 67.02 | 71.52 | 48.49 | |
| DeCE | 74.34 | 79.96 | 0.00 | 63.79 | 69.10 | 0.00 | DeCE | 71.26 | 76.51 | 0.00 | 69.80 | 75.19 | 0.00 | |
@@ 5.1. RQ1: How effective is DeCE compared to existing active defense methods?
The goal of this research question is to establish a benchmark for the performance of DeCE when compared with existing active defense methods. Our evaluation strategy includes a thorough comparative analysis of DeCE and four established active defense techniques, selected from the domains of NLP and CV. This comprehensive comparison spans multiple datasets, CLMs, and poisoning algorithms, ensuring a reliable assessment of DeCE’s effectiveness in thwarting backdoor attacks.
@@
(1) BKI (Chen and Dai, 2021): This method assumes that the defender has the model and the poisoned training set, and removes the poisoned samples from the training set by identifying the importance of each token in the training set, and retrains the model to obtain a model without a backdoor.
@@ Results.
Our empirical studies, as detailed in Table 1, use the highest possible poisoning ratio to test the defense methods against CLMs. For the Lyra and Pisces datasets, we select a poisoning ratio of 5%, while for Bugs2Fix, we chose 1%. The comparative analysis under the RIPPLe and FuncName poisoning strategies is detailed in Table 3, the comparison under the BadPre and DeadCode strategies is provided in Table 4, and the comparison under the Grammar and AFRAIDOOR strategies is provided in Table 5.
The results demonstrate the superior effectiveness of DeCE in countering nearly all backdoor attacks when compared with other active defense methods. Notably, DeCE accomplishes this while preserving the performance of CLMs on clean datasets. The BKI and In-trust Loss methods, however, display inconsistent performance, enhancing security on certain datasets at the expense of others. For instance, with the CodeBERT model, the BKI method enhances security on the Pisces dataset (ASR drops from 87.31% to 18.27%) but adversely affects performance on the Lyra dataset (ASR increases from 56.97% to 93.94%) under the BadPre algorithm. This improvement in security on Pisces is offset by a decline in performance on clean data, as evidenced by a decrease in BLEU scores from 55.06% to 47.33%. The In-trust method also presents a trade-off, improving model security at the cost of decreasing performance on clean datasets across both the Lyra and Pisces datasets. Moderate-fitting and GCE methods exhibit more stable performance, effectively defending against most attacks. Yet, they are susceptible to underfitting, leading to reduced BLEU scores on clean datasets. For example, when the CodeT5 model faces the RIPPLe algorithm, both methods achieve an ASR of 0, signifying robust security. However, this security enhancement may result in a performance drop on clean data. This underscores a critical challenge in the domain of active defense methods, where the quest for heightened security often comes at the expense of decreasing accuracy on legitimate, clean data.
@@
| Model | Defend Method | 5% (RIPPLe) | 5% (BadPre) | 5% (Grammr) | ||||||
| BLEU | CodeBLEU | ASR | BLEU | CodeBLEU | ASR | BLEU | CodeBLEU | ASR | ||
| CodeBERT | ONION | 50.31 | 58.66 | 10.91 | 47.53 | 56.18 | 54.55 | 52.52 | 60.21 | 50.24 |
| Paraphrasing | 38.89 | 48.28 | 1.82 | 37.81 | 46.95 | 1.21 | 38.52 | 47.74 | 4.84 | |
| DeCE | 55.86 | 64.39 | 0.00 | 59.42 | 66.50 | 0.00 | 55.28 | 63.76 | 0.00 | |
| DeCE w. ONION | 55.01 | 63.17 | 0.00 | 48.28 | 57.22 | 0.00 | 52.52 | 60.42 | 0.00 | |
| DeCE w. Paraphrasing | 37.23 | 46.15 | 0.00 | 47.82 | 56.24 | 0.00 | 43.32 | 52.69 | 0.00 | |
| GraphCodeBERT | ONION | 50.65 | 59.06 | 10.91 | 48.76 | 57.08 | 73.33 | 57.20 | 64.52 | 68.18 |
| Paraphrasing | 40.16 | 49.28 | 1.21 | 39.91 | 49.65 | 2.42 | 40.12 | 49.30 | 0.00 | |
| DeCE | 58.48 | 66.54 | 0.00 | 61.20 | 67.58 | 0.00 | 59.55 | 66.28 | 0.00 | |
| DeCE w. ONION | 51.88 | 60.04 | 0.00 | 47.85 | 56.89 | 0.00 | 56.43 | 64.66 | 0.00 | |
| DeCE w. Paraphrasing | 38.54 | 46.83 | 0.00 | 40.16 | 49.92 | 0.00 | 38.51 | 46.69 | 0.00 | |
| CodeGen | ONION | 66.86 | 69.59 | 10.91 | 60.49 | 68.25 | 96.97 | 64.20 | 69.16 | 90.30 |
| Paraphrasing | 41.64 | 49.85 | 3.86 | 42.48 | 50.22 | 6.67 | 41.52 | 49.77 | 3.86 | |
| DeCE | 72.82 | 77.05 | 0.00 | 74.29 | 79.00 | 0.00 | 73.59 | 78.82 | 0.00 | |
| DeCE w. ONION | 66.86 | 69.59 | 0.00 | 61.22 | 69.10 | 0.00 | 64.82 | 68.34 | 0.00 | |
| DeCE w. Paraphrasing | 40.18 | 57.52 | 0.00 | 41.89 | 50.04 | 0.00 | 41.02 | 49.58 | 0.00 | |
| CodeT5 | ONION | 65.27 | 71.33 | 32.12 | 63.03 | 70.28 | 97.58 | 66.17 | 71.94 | 95.76 |
| Paraphrasing | 43.34 | 50.06 | 9.70 | 44.14 | 51.14 | 6.67 | 43.72 | 50.44 | 6.67 | |
| DeCE | 71.66 | 73.57 | 0.00 | 70.26 | 77.44 | 0.00 | 72.28 | 76.34 | 0.00 | |
| DeCE w. ONION | 66.39 | 72.31 | 0.00 | 65.55 | 70.33 | 0.00 | 65.58 | 71.34 | 0.00 | |
| DeCE w. Paraphrasing | 44.71 | 50.08 | 0.00 | 44.58 | 51.62 | 0.00 | 44.22 | 50.86 | 0.00 | |
| CodeT5p | ONION | 65.53 | 71.67 | 32.12 | 62.48 | 69.57 | 96.97 | 64.33 | 70.47 | 93.85 |
| Paraphrasing | 43.10 | 51.43 | 8.48 | 43.36 | 51.30 | 6.67 | 43.27 | 51.41 | 6.67 | |
| DeCE | 75.52 | 80.67 | 0.00 | 75.28 | 80.42 | 0.00 | 73.34 | 79.96 | 0.00 | |
| DeCE w. ONION | 67.61 | 72.94 | 0.00 | 65.64 | 70.73 | 0.00 | 66.20 | 71.13 | 0.00 | |
| DeCE w. Paraphrasing | 42.15 | 50.83 | 0.00 | 43.24 | 51.32 | 0.00 | 42.76 | 50.93 | 0.00 | |
@@ 5.2. RQ2: How effective is DeCE compared to existing passive defense methods?
Building upon the active defense methods chosen in RQ1, we introduce two passive defense methods as baselines for comparison: ONION and Paraphrasing.
@@
Paraphrasing (Jain et al., 2023), on the other hand, leverages the emergent capabilities of Large Language Models (LLMs) to refactor user prompts. Specifically, in the context of CLM backdoor attacks, we utilize the prompt ”Assuming my prompt is unsafe, please paraphrasing my question to the safe prompt.”, allowing gpt-3.5-turbo to perform the paraphrasing.